OpenStudy (donnie1999):

------------WILL FAN AND MEDAL----------- At maximum speed, an airplane travels 1720 miles against the wind in 5 hours. Flying with the wind, the plane can travel the same distance in 4 hours. Let x be the maximum speed of the plane and y be the speed of the wind. What is the speed of the plane with no wind?

OpenStudy (donnie1999):

@pinkbubbles

OpenStudy (donnie1999):

There are no answer choices this time :(

OpenStudy (anonymous):

1720/x=5 5x=1720 x=1720/5 x=344 mph against the wind. 1720/y=4 4y=1720 y=1720/4 y=430 mph. with the wind. (430-344)/2 86/2=43 mph. is the speed of the wind. 344+43=387 430-43=387 mph. speed of the plane in still air.

OpenStudy (donnie1999):

so the answer is 387?

OpenStudy (anonymous):

Yup! :)